OpenStudy (anonymous):

An athlete completes one round of a circular track of radius R in 40 seconds.What will be his displacement at the end of 2 min and 20sec?

OpenStudy (anonymous):

so here is what I did, since in 40 sec, it travelled 2(pi)R, so in 140 sec, it moved 7(pi) R, but this is wrong?

OpenStudy (anonymous):

the answer I have is 2R? how?

OpenStudy (shayaan_mustafa):

Yes I think so it should be 2R.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[(pi \times R ^{2})\div(140)\]

OpenStudy (anonymous):

Excuse me, \[(pi \times R^{2}) \div (140/40)\]

OpenStudy (shayaan_mustafa):

First find velocity from the given data. use the equation\[\Large v=\frac{s}{t}\]

OpenStudy (shayaan_mustafa):

\[\Large s=2pir\] t=40s

OpenStudy (shayaan_mustafa):

yes. substitute your velocity in again that formula for 2min and 40sec calculation.

OpenStudy (anonymous):

so 2piR/40 x 140= 7 pi R

OpenStudy (shayaan_mustafa):

??? how? wrong. 7piR. ?? it must be 2piR

OpenStudy (anonymous):

circ = 2piR you're right so, 2piR(140/40) I always get circ and area mixed up

OpenStudy (shayaan_mustafa):

does this calculation giving you 7piR??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

circ = 2piR you're right so, 2piR(140/40) I always get circ and area mixed up

OpenStudy (anonymous):

circ = 2piR you're right so, 2piR(140/40) I always get circ and area mixed up

OpenStudy (anonymous):

so is the answer 7 pi R? or 2R?

OpenStudy (shayaan_mustafa):

my calculations are giving me 2piR

OpenStudy (anonymous):

2piR/40 x 140 this was your equation?

OpenStudy (shayaan_mustafa):

\[v=\frac{s}{t}\]\[v=\frac{2piR}{40}\]\[v=\frac{piR}{20}\]Now use this \[s=v*t\]\[s=\frac{piR}{\cancel{20}}*\cancel{40}2\]\[s=2piR\]This is how I am trying to do this dear.

OpenStudy (anonymous):

since velocity is the same so shouldn't we put 140 sec instead of 40 sec in the s=vxt equation?

OpenStudy (anonymous):

2piR(140/40)=7piR If you take off all but the fraction of the last lap, you get: piR voila

OpenStudy (anonymous):

displacement = piR

OpenStudy (shayaan_mustafa):

oh... it is 2min and 20 sec. why i thought it is 80 sec. yes you will put 140s

OpenStudy (anonymous):

displacement = piR where 40s is the time for 1 lap, 140s is the time of full run R = radius of circular lap

OpenStudy (shayaan_mustafa):

displacement = piR. how? not halfway track. it is full track though.. isn't it?

OpenStudy (anonymous):

DISPLACEMENT = distance from START

OpenStudy (shayaan_mustafa):

the displacemen for half is + and for next half it is -. you mean this? inclusive directions?

OpenStudy (anonymous):

140s/40s = 3.5 laps .5 = displacement .5*2piR = piR

OpenStudy (shayaan_mustafa):

@physicsme are you getting this point?

OpenStudy (anonymous):

I think you might be confusing Displacement with total distance...@physicsme

OpenStudy (shayaan_mustafa):

Turning test making abuses.

OpenStudy (anonymous):

gimme a medal or something, guys

OpenStudy (anonymous):

yeah, I got it, hit me right after the computer shut down, @Dyiliq you're right, i was mixing displacement with distance. thanks @Shayaan_Mustafa and @Dyiliq :)